CAT Chat Minutes

January 12, 2004

 

General
The viewgraph of long range operations schedule was projected.   The high current run scheduled for February 2004, has been scratched until thermal analyses are completed.   Until Mohan Ramanathan's documentation is complete, we will be running at the normal level (100mA).   Two kicker ceramics are being changed out this shutdown and improvements will be discussed at TWG by Sushil Sharma next week.   There are three other kickers that will be changed out at a later date.  

For the recent 225 mA studies, there were some RF instabilities but they were able to get around them.   The dosimetry devices on the insertion devices were all changed out before the study to determine the doses incurred from this study.   There were no vacuum problems noted during the high mode run.  

Machine Stats - R. Klaffky:   The year ended with a 136-hour fill that improved availability to 97.4% and mean time between faults to 56 hours.

Faults:   During the last week of operations, the FEEPS on Sector 34ID was reset after a water-flow adjustment, and the FEEPS on 14- and 31-ID also required resetting.

Beamline Downtime Report K. Beyer:   Kevin Beyer, AOD floor coordinator, presented the Beamline Downtime Report and explained some of the changes that will be made .   At present, you can search by Sector or Group.   The APS will put this on the web so that users can check downtime reports on their beamlines .

Bill Ruzicka talked about the long-term plan for 300 milliamps and its possible effect on optics, and stated that the APS will provide technical assistance for higher current operation.

Floor Operations - B. Wesolowski :   Bill introduced the new floor coordinator, Rick Spence, Building 438.   Bill presented the new restart packages for the cryo - pumps which need to be re-fabricated.

Safety E. Chang:   January inspections for Building 431 have been completed.  Inspections are scheduled for buildings 432 and 433.